An oil film of thickness 0.12 micrometer rests on a pool of water. If a ray strikes the film at an angle of 60°,what is the wavelength reflected in the first order?
1
Expert's answer
2020-09-10T14:40:03-0400

The condition of the constructive interference in the oil-water film (see https://en.wikipedia.org/wiki/Thin-film_interference#Oil_film):


"2d\sqrt{n_{oil}^2 - \sin^2i} + \dfrac{\lambda}{2} = m\lambda"

where "n_{oil} = 1.5" is the index of refraction of oil, "i = 60\degree" is the incident angle, "d = 0.12\times10^{-6}m" is the thickness of the film and "m = 1,2,3,...". The first wavelenght will be at "m = 1", thus:


"\lambda = 4d\sqrt{n_{oil}^2 - \sin^2i} = 4\cdot 0.12\times 10^{-6}\sqrt{1.5^2 - \dfrac{3}{4}} \approx 623.5\space nm"

Answer. 623.5 nm.